OpenStudy (anonymous):

(7/10) to the 3rd power

OpenStudy (anonymous):

0.343

OpenStudy (anonymous):

How'd u get that?

OpenStudy (anonymous):

it;s supposed to be a simplified fraction

OpenStudy (anonymous):

7^3/10^3 = 343/1000

OpenStudy (anonymous):

Scientific Calculator. :)

OpenStudy (unklerhaukus):

\[({7 \over 10 })^3={7 \over 10}.{7 \over 10}.{7 \over 10}={343 \over1000}=0.343\]